OpenStudy (anonymous):

Let me try again... (a-4)/(-a^3+3a^2) multiplied by (a^2-2a-3)

OpenStudy (anonymous):

factor the denominator as a^2(-a+3)=-a^2(a-3) factor the multiplier as (a-3)(a+1) now (a-4)(a-3)(a+1)/-a^2(a-3) cancel a-3 -(a-4)(a+1)/a^2 can also be written (4-a)(a+1)/a^2

OpenStudy (anonymous):

(a-4)/(-a^3+3a^2) = (a-4) / -a² ( a-3) a² - 2a - 3 = ( a +1) ( a-3) Multiply: => (a-4) ( a +1) ( a-3) / -a² (a-3) = - (a-4) ( a +1)/ a²